Operations Report – January 21st

Operations Report

Subject Line: Crew 187 Operations Report 21JAN2018

Crew 187 Operations Report 21012018

SOL: 09

Name of person filing report: Luis Jose Diaz Lopez

Non-nominal systems: Water heater.

Notes on non-nominal systems: Generator system limping along with a now-nominal routine.

Generator (hours run): 15h 10min

Generator turned off, charging battery at 10h:58m

Generator turned on at 19h:00m

Solar— SOC

@ 08:42: 65%

@ 18:11: 69%

Diesel: It can’t be measure. Explained in the Report 19JAN2018

Propane: 95%

Ethanol Free Gasoline (5 Gallon containers for ATV): 4.75 Gallons

Water (trailer): 0 Gallons

Water (static): 230 Gallons

Trailer to Static Pump used: Yes

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 01300547 Gallons

Toilet tank emptied: No

ATVs Used: None

Oil Added? No

ATV Fuel Used: – Gallons

# Hours the ATVs were used today: –

Notes on ATVs: ATVs were nominal.

Deimos rover used: No

Hours: 101.8 h0

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

Sojourner rover used: Assigned to director only.

Hours: Director discretional hours

Beginning charge:

Ending charge: –

Currently charging: Maybe

Spirit rover used: Yes

Hours: 12.7h

Beginning charge: 100%

Ending charge: –

Currently charging: No

Opportunity rover used: Yes

Hours: 5.6h

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Curiosity rover used: No

Hours: 39h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

HabCar used and why, where? No

General notes and comments: The power generator was running for several hours because there were many clouds in the morning that did not allow the solar panels to charge the batteries. Water heater was disconnected because it’s not working.

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal. Today we got a refill of 125 gallons of water from the water (trailer).

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: None

Sol Summary – January 21st

Crew 187 Sol 9 Summary Report 21JAN2018

Sol 9

Summary Title: Zombie

Author’s name: Cynthia Fuertes Panizo

Mission Status: All nominal

Sol Activity Summary:

Is there someone in Earth?

This day we decided not to do the exercises and start the day late, so at 8:15 we ate breakfast and every Martian started to work in their project, our alarm was delayed an hour. The tiredness begins to be noticed because the naps are more common between meals. At 11:00 our Martian chef, Danton, started to cooked a delicious food: Chicken with cheese, pasta and vegetables. We ate together before the expedition began. At 11:30, Tania, David, Luis and Oscar started the donning time. They took Opportunity and Spirit for this adventure… It’s supposed that they were going to work in El Dorado Canyon, but it was a confusion with the location so they visited a near canyon. However, they took some samples of the halophiles needed. The path that they follow was bent. The canyon that the Martians explorers visited combines a wide variety of landscapes, from rocks up to 5 meters high, which route land for an approximate of 500 meters between ascents and descents and with different types of soil and frozen water remains.

By 15:00, our chef Danton started to cooked a second lunch for our Martian explorers, a delicious soup and scrambled eggs with chili. They arrived at 15:20 to our home. And 25 minutes later we started to eat together… again!

In the afternoon, we got a refill of 125 gallons from the water (trailer). Now we have water for the rest of our lives in Mars, or at least, we hope so.

David is replanting 12 quinua’s plants on the lysimeter for his project. Meanwhile, I get an advanced programing my augmented reality mobile application. And, as always, Tania is documenting everything for her project.

For dinner, Atila is trying to cook a surprise Martian dinner…. Everybody is waiting to taste it!

Ad Astra,

Cynthia Fuertes Panizo

Commander Crew 187 – MDRS

Look Ahead Plan:

Another EVA is going to take place. The lucky ones are Luis, Atila, David and Danton. They are going to Copernicus Hwy.

Anomalies in work:


Weather: Temperature: -1°C, Fair weather, Winds speed: NNE 8 KPH, Humidity 43%, Barometer 30.22 in

Crew Physical Status: Healthy

Reports to be filed: Sol Summary, GeenHab Report, Operations Report, EVA Request, EVA Report and HSO Report.

Support Requested:

Please sent us distillated water.

GreenHab Report – January 21st

GreenHab Report

Hernán David Mateus Jiménez


Environmental control:

Shade cloth on

Working Hour: 18:20
Inside temp at working hour: 19° C
Outside temp during working hours: -1° C
Inside temperature high: 24° C
Inside temperature low: 15° C
Inside humidity: 48 %RH

Inside humidity high: 48 %RH
Inside humidity low: 16 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: None

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 18:30

Changes to research plants: none

Aquaponics: Atila is assembling it. He is waiting for the amount of water to fill the tank, this is necessary to do the test.

Narrative: The lysimeter is taking data. Today 12 quinoa plants were replanted in the lysimeter over Martian soil. Atila and I have made some experiments to discovers what is a good proportion between soil and water to get the best texture, we have figured out that you need 500 mL of water for 2800 g of soil.

EVA Report – January 21st

Purpose of EVA: Primary: Mid-range testing of the Cóndor Space Suit Simulator, Mid-range EVA protocol testing, Sample collection for possible Halophile identification, further exploration of El Dorado Canyon.

Secondary: Outreach and Educational graphic material collection. Relevant sample collection.

Participants: Oscar, Luis, David, and Tania.

Narrative: Today marked the last of the mid-range EVAs programmed, leading to longer and even more interesting exploration. The weather was favorable, but windy. The crew donned the suits and exited the airlock at 12:30, the four crewmembers then boarded the Rovers Spirit and Opportunity and followed the path planned, towards El Dorado canyon. After approximately 25 minutes of traverse, and after passing Robert’s Rock Garden, the Rovers were parked on the side of the road, and the expedition started walking with Southeast course, towards the entrance of the canyon, upon finishing the rocky plains, two crewmembers had to return to the Rover, given that the aerial drone was left behind. Upon reuniting, the crew resumed the traverse until the entrance of the canyon. More herbivore and cougar prints were found in the area, but not as fresh as on the previous EVA to the canyon. The crew entered the canyon and followed the dry riverbed. Along this path, water samples were taken, aerial footage with the drone, as well as traditional photography. The crew followed the path until the eastern exit of the canyon, heading south to meet Cow Dung road again. However, the sampling sites that were found during the firs EVA to the area were not found, since they probably belong to a subsidiary structure, south of El Dorado canyon. It is suggested that further detailed mapping of this area is undertaken, to identify all the morphologies present. While returning through Cow Dung road, heading north, a crewmember’s cell phone (used as a compass and camera) was lost, so a search formation was done, and the crew headed south, being able to find the device, and resuming the walk back north. The walk lasted about 25 minutes, reuniting with the Rovers and heading back to the habitat. The EVA ended at 15:20 with the ingress of the crew from the airlock to the habitat.

The aerial footage is being processed, as well as the performance of the Cóndor Space Suit Simulator, which will be subjected to detailed inspection tomorrow.

The only incident related with the EVA was the lost and recovery of the cell phone.

Copyright © The Mars Society. All rights reserved. | Main Site